# Efficient Bayesian analysis of kilonovae and GRB afterglows with fiesta

### v1

@@ 3 Surrogate training
Table 1: Overview of fiesta surrogates for different physical GRB afterglow models (afterglowpy and pyblastafterglow) and the KN model (possis). We list the parameters, their ranges in the training data set, as well as the architectures employed for the ML surrogates. The models marked in bold are the best-performing surrogate architectures and are used in Secs. 4 and 5. The bar in the kilonova parameters indicates mass-averaged quantities.
@@
jet core angle [rad] $\theta_{\text{c}}$ $[0.01,\pi/5]$
@@
log interstellar medium density [cm-3] $\log_{10}(n_{\text{ism}})$ $[-6,2]$
electron spectrum power index $p$ $[2,3]$
log electron energy fraction $\log_{10}(\epsilon_{e})$ $[-4,0]$
log magnetic energy fraction $\log_{10}(\epsilon_{B})$ $[-8,0]$
@@
inclination [rad] $\iota$ $[0,\pi/2]$
log dynamical ejecta mass [M] $\log_{10}(m_{\text{ej, dyn}})$ $[-3,-1.3]$
@@
log wind ejecta mass [M] $\log_{10}(m_{\text{ej, wind}})$ $[-2,-0.9]$ LightcurveModel (MLP)
@@
To evaluate the likelihood function in Eq. (2) efficiently, fiesta provides surrogates that determine the expected magnitudes $m^{\star}(t_{j},\@vec{\!\theta}\,)$ for a given parameter point ${\theta}$ through ML surrogates.
For KN models, previous works have demonstrated the capability of ML techniques to replace the expensive lightcurve generation of radiative transfer models, which allowed their use in stochastic samplers (Almualla et al., 2021; Ristic et al., 2022; Kedia et al., 2023; Lukošiūtė et al., 2022; Ristic et al., 2023; Ford et al., 2024; Saha et al., 2024; King et al., 2025). For GRB afterglows, many fast semi-analytic models are available, such that the use of ML surrogates has not been as essential, though an increasing number of recent works have introduced ML techniques to this area. In Lin et al. (2021), the authors linearly interpolate a fixed table of GRB afterglow lightcurves from the prescription in Lamb et al. (2018) to accelerate the likelihood evaluation, though the interstellar medium density and other microphysical parameters were kept fixed. A surrogate model for the X-ray emission in afterglowpy was trained in Sarin et al. (2021) to analyze the Chandra transient CDF-S XT1. In Boersma & van Leeuwen (2022), DeepGlow was introduced, a python package that emulates boxfit (van Eerten et al., 2012) lightcurves through a neural network (NN). In Wallace & Sarin (2024), the authors trained an NN for the afterglow model of Lamb et al. (2018). Rinaldi et al. (2024) suggested developing an ML surrogate based on the afterglow model developed in Warren et al. (2022), but postponed the implementation to future work. In Aksulu et al. (2020, 2022) the authors used Gaussian processes to model the likelihood function in Bayesian analysis, though the expected lightcurves were computed directly with scalefit (Ryan et al., 2015).
In the present work, we introduce the first large-scale surrogates for the state-of-the-art afterglow models afterglowpy and pyblastafterglow, covering the radio to hard X-ray emission over a timespan of $10^{-4}-2\times 10^{3}$ days. Additionally, we extend the work of Almualla et al. (2021); Anand et al. (2023) and train a new KN surrogate with an updated version of the possis code. Our surrogates provide a large speed-up, since the possis Monte-Carlo radiative transfer code (Bulla, 2019, 2023) takes on the order of $\sim 1000$ CPU-hours to predict a lightcurve. Likewise, depending on the settings, a GRB afterglow simulation in afterglowpy takes on the order of 0.1 - 10 seconds, while for pyblastafterglow the computation time may exceed several minutes. In the following subsections, we provide details about the implemented ML approaches in fiesta and present the surrogates we obtained for the GRB afterglow and KN models.
@@ 3.1 Surrogate types and architectures
To create a surrogate for fiesta, an NN learns the relationship between the input ${\theta}$, i.e., the parameters of the model, and the output ${y}$, i.e., the spectral or photometric flux. The surrogate models thus interpolate a precomputed training data set that consists of many evaluations of the physical base model, e.g., possis or pyblastafterglow, on various combinations of input parameters.
Specifically, the output ${y}$ could either represent the integrated photometric fluxes (or bandpass magnitudes) in predefined frequency filters, or yield the spectral flux density $F_{\nu}$ across a continuum of frequencies. In previous works, surrogate models typically predicted magnitudes for a set of predefined filters with fixed wavelength (Almualla et al., 2021; Pang et al., 2023; Peng et al., 2024). Training on the spectral flux densities provides more flexibility, as the surrogate does not need to be retrained if a new filter becomes available, and the surrogate’s output can be further processed to account for arbitrary redshifts or extinction effects. Surrogates trained on $F_{\nu}$ will thus return a 2D-array containing the fluxes across time along one dimension, and across frequency along the other dimension. fiesta implements both types of surrogates, which we refer to as FluxModel class for the latter kind of surrogate, whereas the surrogates that are trained in the traditional approach on bandpass magnitudes are referred to as the LightcurveModel class.
Independent of the particular type of surrogate model, fiesta employs two kinds of NN architectures, namely the simple feed-forward NN and conditional variational autoencoders (cVAEs). The feed-forward NN, in this work containing three hidden layers, is used to train the relationship between the input parameters ${\theta}$ and the coefficients ${c}$ of the principal component analysis (PCA) of the training data. The training simply minimizes the mean squared error on the PCA coefficients as loss function
@@
where ${\phi}$ are the NN weights, ${c}^{\text{ (train)}}_{j}$ are the PCA coefficients of the training data ${y}$, and ${c}^{\text{ (predict)}}_{j}$ are the coefficients the NN would predict. To then later infer the bandpass magnitude or flux density ${y}$, the inverse PCA decomposition is applied to ${c}$. The feed-forward architecture is thus used both for LightcurveModel and FluxModel surrogates.
The other implemented architecture is the cVAE (Kingma & Welling, 2013; Rezende et al., 2014). This approach is inspired by Lukošiūtė et al. (2022), where cVAEs were used to predict KN spectra. Unlike their work, our setup predicts the flux density across a fixed grid of times and frequencies, and thus time is not a training parameter. Moreover, we extend this approach to GRB afterglows.
In the cVAE architecture, an encoder and decoder are trained simultaneously on the spectral flux densities directly. The encoder takes the parameters ${\theta}$ and the flux ${y}$ as inputs and maps them to the latent parameters ${\mu}$ and ${\sigma}$. These parameters represent the variational distribution of the latent space from which the decoder reconstructs ${y}$. Specifically, the latent vector is drawn according to ${z}\sim\mathcal{N}(\@vec{\mu},\,\text{diag}(\@vec{\sigma}))$ and serves as input to the decoder, together with ${\theta}$. The encoder and decoder are trained on minimizing the joint loss function
$\begin{split}&L(\@vec{\phi}_{\text{encoder}},\@vec{\phi}_{\text{decoder}})=\\ &\frac{1}{n_{\text{train}}}\sum_{j=1}^{n_{\text{train}}}\bigg{(}\frac{1}{2}\left(1+\ln(\@vec{\sigma}_{j})^{2}-\@vec{\sigma}_{j}^{2}-\@vec{\mu}_{j}^{2}\right)+\lVert\@vec{y}^{\text{ (train)}}_{j}-\@vec{y}^{\text{ (predict)}}_{j}\rVert^{2}\bigg{)}\,,\end{split}$ (5)
where ${\phi}$ represent the NN weights, ${\sigma}_{j}$, ${\mu}_{j}$, and ${y}^{\text{ (predict)}}_{j}$ the predicted parameters for the variational distribution and the flux, and ${y}^{\text{ (train)}}_{j}$ the training data. The first term in Eq. (5) represents the Kullback-Leibler divergence between the variational distribution and the standard normal distribution, while the second term is the reconstruction loss. Since the decoder is conditioned with the parameters ${\theta}$, its output after training will not depend much on the latent vector, and for the actual flux prediction we set ${z}=(0,\dots,0)$. Because the cVAE is computationally more expensive to train, it was only implemented for FluxModel surrogates, where just a single surrogate is trained, in contrast to the LightcurveModel where each photometric filter constitutes its own surrogate.
All NNs are implemented through the Flax API (Heek et al., 2024).
@@ 3.2 GRB afterglow surrogates
Figure 1: Benchmarks of two surrogates for the afterglowpy Gaussian jet model. We show the error distributions of the surrogate predictions against a test data set of size $n_{\text{test}}=7500$. The different panels show the error of the predicted photometric lightcurve in the different bandpasses. The left panels show the distribution of the mean squared error across the entire lightcurve, i.e. the average deviation between $10^{-4}$$2000$ days. The right panels show the mismatch distribution across the test data set of between predicted magnitude and test data as a function of time. The figure compares two different surrogates, one using the MLP architecture (blue), the other a cVAE (green).
@@
Figure 2: Benchmarks of the surrogates for the pyblastafterglow Gaussian jet model. We show the deviations of surrogate predictions against a test data set of size $n_{\text{test}}=7232$. Figure layout as in Fig. 1.
@@
In fiesta, we include surrogates for the GRB afterglow models afterglowpy (Ryan et al., 2020) and pyblastafterglow (Nedora et al., 2024). Table 1 shows the parameter ranges and trained architectures for these models. We note that these surrogates are specific to a structured Gaussian jet. While we also created GRB afterglow surrogates for the tophat jet models, these surrogates are of less relevance in real-life applications (Ryan et al., 2020; Salafia & Ghirlanda, 2022) and we discuss their performance in Appendix A. We simply note here that the tophat jet surrogates generally outperform the surrogates for the Gaussian jet in terms of accuracy with respect to the ground truth from the base model, due to their simpler physical behavior.
@@
Both afterglowpy and pyblastafterglow assume that the jet of the GRB can be modeled as a relativistic fluid shell that propagates trough a cold, ambient interstellar medium with constant density $n_{\rm ism}$ as a shockwave. The shock jumping conditions can then be used to determine the shell dynamics analytically given the initial kinetic energy $E_{0}$ and bulk Lorentz factor $\Gamma_{0}$ (Nava et al., 2013; Ryan et al., 2020). Both afterglowpy and the pyblastafterglow runs in the present work ignore reverse shock contributions and only include the forward shock. If the jet has a Gaussian structure, then the jet is evolved as a collection of individual, independent, annular shells each of them assigned an energy according to
@@
$E(\theta)=\begin{cases}E_{0}\exp\left(-\frac{\theta^{2}}{2\theta_{c}^{2}}\right)\qquad\text{if $\theta\leq\theta_{\rm w}$}\\ 0\qquad\qquad\qquad\qquad\text{else}\end{cases}\ ,$ (6)
@@
where $\theta$ is the polar angle from the jet’s center axis and $\theta_{\rm c}$ the core angle parameter of the jet. The wing angle $\theta_{\rm w}$ is parameterized in Table 1 through the factor $\alpha_{\text{w}}=\frac{\theta_{\rm w}}{\theta_{\rm c}}$. The radiation is then determined as synchrotron emission from the accelerated electrons behind the shock front that receive a fraction $\epsilon_{e}$ of the shock energy as acceleration. The magnetic field in the downstream shock is given through a fraction $\epsilon_{B}$ of the shock energy. While employing a similar semi-analytical framework, afterglowpy and pyblastafterglow have notable differences in their specific implementation. Most notably, afterglowpy assumes a fixed electron spectrum $N_{e}(\gamma_{e})\propto\gamma_{e}^{-p}$ with the electron spectrum power index $p$, whereas pyblastafterglow numerically evolves the electron distribution and only the injected electron distribution follows this power law. Further differences regard the implementation of the jet spreading and initial coasting phase and the methods used to compute the synchrotron radiation spectra.
@@
For both GRB afterglow models, we train two surrogates of the FluxModel type, one with an MLP architecture, the other with the cVAE architecture. Each surrogate was trained with input parameters as listed in Table 1. The training data was randomly drawn from the ranges specified there. The training data set for the afterglowpy Gaussian jet surrogate encompasses $n_{\text{train}}=80000$ flux density calculations, the set for the pyblastafterglow surrogate $n_{\text{train}}=91670$. The surrogates were trained on standardized $\ln(F_{\nu})$, and standardized parameter samples ${\theta}$. After different attempts with similar results, the number of PCA coefficients for the MLP training was set to 50 and the cVAE trained on a down-sampled flux density array of size $42\times 57$. On a H100 GPU, training took about 2.2 h for the cVAE and 0.3 h for the MLP architecture.
@@
In Fig. 1, we show the performance of the surrogates for the afterglowpy Gaussian jet model when trying to predict the lightcurves of a test dataset, i.e., a dataset that was not used for training. The overall square error is confined to $\lesssim 0.1$ mag across different bandpasses and does not vary notably between the MLP or cVAE architectures. The panels on the right hand-side show the distribution of the absolute mismatch over time. While this mismatch is mostly within 0.3 mag, some outlier predictions show stronger deviations from the test data. Specifically, for the cVAE architecture, $\sim$6of all test data samples exceed a mismatch of 1 mag at some point along the test lightcurve.
@@
In Fig. 2, we show the performance of the corresponding pyblastafterglow surrogates for the Gaussian jet. The error is typically higher compared to the surrogate for afterglowpy, which we attribute to higher variability in the training data, given the additional features of pyblastafterglow over afterglowpy. This is also represented in the mismatch, which typically falls in the range of $\lesssim 0.4$ mag, though for $\sim$10of the test data samples the mismatch exceeds 1 mag at least once along the lightcurve. We also note that the X-ray filter in the bottom right panel shows somewhat higher mismatches from the predictions. This is due to the fact that we cropped the training data below $2\cdot 10^{-22}$ mJys at 10 pc (corresponding roughly to the 70th absolute magnitude) due to numerical noise in the pyblastafterglow flux computation at very low brightness. This cutoff mostly affects the high energy emission, and since the surrogates struggle to reproduce this hard cut, the typical surrogate mismatch in this channel is higher than for the lower frequency filters. As this concerns only lightcurves far below the detection limit, this is no issue in real-life applications.
@@
The MLP and cVAE architectures perform very similarly for the GRB afterglow surrogates, though the cVAE tends to have a slightly smaller absolute mismatch in the lightcurve at later times. This is especially true for the pyblastafterglow model. Overall, the cVAEs also tend to have the smaller average square error as shown by the distributions in the left panels of Fig. 1 and 2. For these reasons, we set the cVAE as the default architecture to be used for the analyses later in Sec. 4 and 5.

### v2

@@ 3 Surrogate training
Table 1: Overview of fiesta surrogates for different models. [^3] We list the parameters, their ranges in the training data set, as well as the architectures employed for the ML surrogates. The models marked in bold are the best-performing surrogate architectures and are used in Secs. 4 and 5. The bar above the kilonova parameters indicates mass-averaged quantities.
@@
jet core angle [rad] $\theta_{\text{c}}$ $0.01,\pi/5]$
@@
log interstellar medium density [cm-3] $\log_{10}(n_{\text{ism}})$ [6,2][-6,2]
electron power index $p$ $[2,3]$
log electron energy fraction $\log_{10}(\epsilon_{e})$ [4,0][-4,0]
log magnetic energy fraction $\log_{10}(\epsilon_{B})$ [8,0][-8,0]
@@
inclination [rad] $\iota$ $0,\pi/2]$
log dynamical ejecta mass [M] $\log_{10}(m_{\text{ej, dyn}})$ [3,1.3][-3,-1.3]
@@
log wind ejecta mass [M] $\log_{10}(m_{\text{ej, wind}})$ [2,0.9][-2,-0.9] LightcurveModel (MLP)
@@
To evaluate the likelihood function in Eq. (3) efficiently, fiesta provides surrogates that determine the expected magnitudes, $m^{\star}(t_{j},\@vec{\!\theta}\,)$, for a given parameter point $\@vec{\theta}$ through ML surrogates. For KN models, previous works have demonstrated the capability of ML techniques to replace the expensive light curve generation of radiative transfer models, which enabled their use in stochastic samplers (Almualla:2021znj; Ristic:2021ksz; Kedia:2022onl; Lukosiute:2022tmd; Ristic:2023ywr; Ford:2023bdi; Saha:2023zse; King:2025tqo). For GRB afterglows, many fast semi-analytic models are available, so the use of ML surrogates has not been as essential; however, an increasing number of recent works have introduced ML techniques to this area. In Lin:2021auf, the authors linearly interpolated a fixed table of GRB afterglow light curves from the prescription in Lamb:2018ohw to accelerate the likelihood evaluation, although the interstellar medium density and other microphysical parameters were kept fixed. A surrogate model for the X-ray emission in afterglowpy was trained in Sarin:2021yya to analyze the Chandra transient CDF-S XT1. In Boersma:2022qtg, DeepGlow was introduced, a python package that emulates boxfit (vanEerten:2011yn) light curves through a neural network (NN). In Wallace:2024pmx, the authors trained an NN for the afterglow model of Lamb:2018ohw. Rinaldi:2024xjz suggested developing an ML surrogate based on the afterglow model developed in Warren:2021whb, but postponed the implementation to future work. In Aksulu:2020hnl; Aksulu:2021crt, the authors used Gaussian processes to model the likelihood function in Bayesian analysis, although the expected light curves were computed directly with scalefit (Ryan:2014nea).
In the present work, we introduce the first large-scale surrogates for the state-of-the-art afterglow models afterglowpy and pyblastafterglow, covering the radio to hard X-ray emission over a timespan of $10^{-4}-2\times 10^{3}$ days. Additionally, we extend the work of Almualla:2021znj; Anand:2023jbz and train a new KN surrogate with an updated version of the possis code. Our surrogates provide a large speed-up, since the possis Monte Carlo radiative transfer code (Bulla:2019muo; Bulla:2022mwo) takes on the order of $\sim 1000$ CPU-hours to predict a light curve. Likewise, depending on the settings, a GRB afterglow simulation in afterglowpy takes on the order of 0.1 - 10 seconds; whereas for pyblastafterglow, the computation time may exceed several minutes. In the following subsections, we provide details about the implemented ML approaches in fiesta and present the surrogates we obtained for the GRB afterglow and KN models.
@@ 3.1 Surrogate types and architectures
To create a surrogate for fiesta, a NN learns the relationship between the input $\@vec{\theta}$, i.e., the parameters of the model, and the output y\@vec{y} (i.e., the flux). The surrogate models thus interpolate a precomputed training data set that consists of many evaluations of the physical base model (e.g., possis or pyblastafterglow) on various combinations of input parameters.
Specifically, the output y\@vec{y} could either represent the magnitudes in predefined frequency filters, or yield the spectral flux density $F_{\nu}$ across a continuum of frequencies. In previous works, surrogate models typically predicted magnitudes for a set of predefined filters with fixed wavelength (Almualla:2021znj; Pang:2022rzc; Peng:2024jqe). Training on the spectral flux densities provides more flexibility, as the surrogate does not need to be retrained if a new filter becomes available, and the surrogate’s output can be further processed to account for arbitrary redshifts or extinction effects. Surrogates trained on $F_{\nu}$ will thus return a 2D-array containing the flux density across time along one dimension, and across frequency along the other dimension. fiesta implements both types of surrogates, which we refer to as FluxModel class for the latter kind of surrogate, whereas the surrogates that are trained in the traditional approach on passband magnitudes are referred to as the LightcurveModel class.
Regardless of the particular type of surrogate model, fiesta employs two kinds of NN architectures, namely the simple feed-forward multilayer perceptron (MLP) and the conditional variational autoencoder (cVAE). The feed-forward NN, in this work containing three hidden layers, is used to train the relationship between the input parameters $\@vec{\theta}$ and the coefficients c\@vec{c} of the principal component analysis (PCA) of the training data. The training simply minimizes the mean squared error on the PCA coefficients as loss function,
@@
where $\@vec{\phi}$ are the NN weights, $\@vec{c}^{\text{ (train)}}_{j}$ are the PCA coefficients of the training data y\@vec{y}, and $\@vec{c}^{\text{ (predict)}}_{j}$ are the coefficients the NN would predict. The passband magnitude or flux density y\@vec{y} is then determined by applying the inverse PCA decomposition to c\@vec{c}. The feed-forward architecture is thus used both for LightcurveModel and FluxModel surrogates.
The other implemented architecture is the cVAE (Kingma:2013hel; Rezende:2014azh). This approach is inspired by Lukosiute:2022tmd, where cVAEs were used to predict KN spectra. In contrast to their work, our setup predicts the flux density across a fixed grid of times and frequencies, and thus time is not a training parameter. Moreover, we extended this approach to GRB afterglows.
In the cVAE architecture, an encoder and decoder are trained simultaneously on the spectral flux densities directly. The encoder takes the parameters $\@vec{\theta}$ and the flux y\@vec{y} as inputs and maps them to the latent parameters $\@vec{\mu}$ and $\@vec{\sigma}$. These parameters represent the variational distribution of the latent space from which the decoder reconstructs y\@vec{y}. Specifically, the latent vector is drawn according to $\@vec{z}\sim\mathcal{N}(\@vec{\mu},\,\text{diag}(\@vec{\sigma}))$ and serves as input to the decoder, together with $\@vec{\theta}$. The encoder and decoder are trained on minimizing the joint loss function,
[…]
where $\@vec{\phi}$ represents the NN weights, $\@vec{\sigma}_{j}$, $\@vec{\mu}_{j}$, and $\@vec{y}^{\text{ (predict)}}_{j}$ are the predicted parameters for the variational distribution and the flux, and $\@vec{y}^{\text{ (train)}}_{j}$ is the training data. The first term in Eq. (7) represents the Kullback-Leibler divergence between the variational distribution and the standard normal distribution, while the second term is the reconstruction loss. Since the decoder is conditioned with the parameters $\@vec{\theta}$, its output after training will not depend much on the latent vector, and for the actual flux prediction we set $\@vec{z}=(0,\dots,0)$. Since the cVAE is computationally more expensive to train, it was only implemented for FluxModel surrogates, where just a single surrogate is trained, in contrast to the LightcurveModel where each photometric filter constitutes its own surrogate. All NNs were implemented through the Flax API (flax2020github).
@@ 3.2 GRB afterglow surrogates
Figure 1: Benchmarks of the two surrogates for the afterglowpy Gaussian jet model. We show the error distributions of the surrogate predictions against a test data set of size $n_{\text{test}}=7500$. The different rows show the error across different passbands. The left panels show the distribution of the mean squared error as defined in Eq. (9). The right panels show the mismatch distribution across the test data set as defined in Eq. (10). The figure compares two different surrogates: one using the MLP architecture (blue) and the other a cVAE (green).
@@
Figure 2: Benchmarks of the two surrogates for the pyblastafterglow Gaussian jet model. We show the deviations of surrogate predictions against a test data set of size $n_{\text{test}}=7232$. Figure layout is the same as in Fig. 1.
@@
In fiesta, we include surrogates for the GRB afterglow models afterglowpy (Ryan:2019fhz) and pyblastafterglow (Nedora:2024vrv). Table 3 shows the parameter ranges and trained architectures for these models. We note that these surrogates are specific to a structured Gaussian jet. While we also created surrogates for the tophat jet models, their relevance in real-life applications is limited (Ryan:2019fhz; Salafia:2022dkz) and we discuss their performance in Appendix A. We simply note here that the tophat jet surrogates generally outperform the surrogates for the Gaussian jet in terms of accuracy due to their simpler physical behavior.
@@
Both afterglowpy and pyblastafterglow assume that the jet of the GRB can be modeled as a relativistic fluid shell that propagates trough a cold, ambient interstellar medium with constant density $n_{\rm ism}$ as a shockwave. The shock jumping conditions can then be used to determine the shell dynamics analytically given the initial kinetic energy $E_{0}$ and bulk Lorentz factor $\Gamma_{0}$ (Nava:2012hq; Ryan:2019fhz). All afterglowpy and the pyblastafterglow runs in the present work ignore reverse shock contributions and only include the forward shock. If the jet has a Gaussian structure, then the jet is evolved as a collection of individual, independent, annular shells each of them assigned an energy according to
@@
$E(\theta)=\begin{cases}E_{0}\exp\left(-\frac{\theta^{2}}{2\theta_{c}^{2}}\right)\qquad\text{if $\theta\leq\theta_{\rm w}$}\\ 0\qquad\qquad\qquad\qquad\text{otherwise}\end{cases}\ ,$ (8)
@@
where $\theta$ is the polar angle from the jet’s center axis and $\theta_{\rm c}$ the core angle parameter of the jet. The wing angle $\theta_{\rm w}$ is parameterized in Table 3 through the factor $\alpha_{\text{w}}=\frac{\theta_{\rm w}}{\theta_{\rm c}}$. Once the dynamics have been determined, the emission is modeled as synchrotron radiation from the electrons accelerated at the shock front, which receive a fraction $\epsilon_{e}$ of the shock energy. The magnetic field in the downstream shock is given through a fraction $\epsilon_{B}$ of the shock energy. While employing a similar semi-analytical framework, afterglowpy and pyblastafterglow have notable differences in their specific implementation. Most notably, afterglowpy assumes that the electron distribution follows a broken power law in which newly shocked electrons are injected with $N_{e,\text{inj}}(\gamma_{e})\propto\gamma_{e}^{-p}$, where $p$ is the electron power index. pyblastafterglow assumes the same power law for the newly shocked electrons, but instead numerically evolves the existing electron distribution. Further differences regard the implementation of the jet spreading and initial coasting phase and the methods used to compute the synchrotron radiation spectra.
@@
For both GRB afterglow models, we trained two surrogates of the FluxModel type: one with an MLP architecture and the other with the cVAE architecture. Each surrogate was trained with input parameters as listed in Table 3. The training data was randomly drawn from the ranges specified there. The training data set for the afterglowpy Gaussian jet surrogate encompasses $n_{\text{train}}=80000$ flux density calculations, the set for the pyblastafterglow surrogate $n_{\text{train}}=91670$. The surrogates were trained on standardized $\ln(F_{\nu})$, and standardized parameter samples $\@vec{\theta}$. After different attempts with similar results, the number of PCA coefficients for the MLP training was set to 50 and the cVAE trained on a down-sampled flux density array of size $42\times 57$. On a H100 GPU, training took about 2.2 h for the cVAE and 0.3 h for the MLP architecture.
We used two different metrics to compare the light curves $m^{\rm pred}(t)$ predicted by the surrogate against a set of test light curves $m^{\rm test}(t)$. These test light curves are from the physical base model and were not part of the training set. The one metric is the mean squared error MSE,
$\text{MSE}^{2}=\int_{\log(t_{\rm min})}^{\log(t_{\rm max})}\frac{(m^{\rm test}(t)-m^{\rm pred}(t))^{2}}{\log(t_{\rm max})-\log(t_{\rm min})}d\log(t)\ ,$ (9)
and the other is the mismatch MIS,
$\text{MIS}(t)=|m^{\rm test}(t)-m^{\rm pred}(t)|\ .$ (10)
@@
In Fig. 1, we show the performance of the surrogates for the afterglowpy Gaussian jet model when trying to predict the magnitudes in different passbands. Overall, the squared error is confined to $\lesssim 0.1$ mag across different photometric filters and does not vary notably between the MLP or cVAE architectures. The panels on the right hand-side show the distribution of the absolute mismatch over time. While this mismatch is mostly within 0.3 mag, some outlier predictions show stronger deviations from the test data. Specifically, for the cVAE architecture, $\sim$6% of all test data samples exceed a mismatch of 1 mag at some point along the test light curve.
@@
In Fig. 2, we show the performance of the corresponding pyblastafterglow surrogates for the Gaussian jet. The deviation from the test data set is typically larger than for the afterglowpy surrogate, which we attribute to higher variability in the training data arising from the additional features of pyblastafterglow. This is also represented in the mismatch, which typically falls in the range of $\lesssim 0.4$ mag, though for $\sim$10% of the test data samples the mismatch exceeds 1 mag at least once along the light curve. We also note that the X-ray filter in the bottom right panel shows somewhat higher mismatches from the predictions. This is due to the fact that we cropped the training data below $2\cdot 10^{-22}$ mJys at 10 pc (corresponding roughly to the 70th absolute magnitude) due to numerical noise in the pyblastafterglow flux computation at very low brightness. This cutoff mostly affects the high frequencies, and since the surrogates struggle to reproduce this hard cut, the typical mismatch in the X-ray filter is higher than for the lower frequency filters. As this concerns only light curves far below the detection limit, this is no issue in real-life applications.
@@
The MLP and cVAE architectures perform very similarly for the GRB afterglow surrogates, though the cVAEs tend to have a slightly smaller absolute mismatch in the light curve at later times. This is especially true for the pyblastafterglow model. Overall, the cVAEs also tend to have the smaller average square error as shown by the distributions in the left panels of Figs. 1 and 2. For these reasons, we set the cVAE as the default architecture to be used for the analyses later in Sects. 4 and 5.